OpenStudy (anonymous):

Do the algebra to show the solutions to (5/3)x^(2/3) - (4/3)x^(1/3)=0 are x=0, 64/125

hartnn (hartnn):

put x^(1/3) =y

hartnn (hartnn):

then x={2/3} = y^2

hartnn (hartnn):

and u will get a quadratic in y